# Preparation of Schrödinger cat quantum state using parametric down-conversion interaction

### v2

@@ I Introduction
One of the most well known type of the non-Gaussian states is the Schrödinger cat (SC) one. The “mainstream” method of preparation of these states is based in photon subtraction and conditional measurement, see Refs. [20, 21, 22, 23, 24, 25] and the review [26]. The values of SC state amplitudes up to $\alpha\sim 2$ were demonstrated using this method. The recent estimates [27, 28] show that it allows to prepare SC states with the amplitudes $\alpha\gtrsim 4$ and fidelity exceeding 99%.
@@
Unfortunately, this method requires a very strong parametric interaction and small linear losses in the signal mode [36]. By now, these conditions were successfully fulfilled only in microwave cryogenic systems with Josephson junctions [37, 38].
In the works [29, 30], joint unitary evolution of the two modes was analyzed numerically. In particular, in Ref. [29], evolution of the Husimi quasi-probability function of the signal mode was explored. However, in the general case, tracing out the pump mode gives a mixed state of the signal one. Therefore, this approach can not be considered as the optimal one.
@@ III Unconditional evolution
In Fig. 2, the probability distribution for the photon number in the signal mode is shown. Similar to the “even” SC state
@@ IV Conditional preparation of the Schrödinger cat state
Suppose that at some moment of the normalized time $\tau_{\rm opt}$, the photon number $\hat{n}_{p}$ in the pump mode is measured. Suppose also that we are interested only in the events where the value $\hat{n}_{p}=0$ is obtained. The schematic diagram of this procedure is shown in Fig. 3.
Figure 3: The scheme of the conditional preparation of the SC state. Red line – signal mode, blue line – pump mode. SPD – single photon detector.
@@
Figure 5: Wigner function of the output state $|{\psi_{\rm out}}\rangle$ at the optimal time $\tau_{\rm opt}$. $\beta=3$
@@
In Fig. 6, deviation of the resulting fidelity from one is plotted as a function of $\beta$. It does not exceed $3\cdot 10^{-3}$ and most probably results from the accumulated errors of the numerical calculations.
The prepared squeezed SC state $|{\psi_{\rm out\,raw}}\rangle$ could be converted to the non-squeezed one $|{\psi_{\rm out}}\rangle$ using the anti-squeezing operation with the same squeeze factor $\-\xi_{\rm prep}$, see Eq. (17). The resulting average photon number can be calculated as follows (see Eq. (30)):
@@
V Discussion
@@ Acknowledgements.
This work was supported by the Russian Science Foundation (project 20-12-00344). The author would like to thank D. Chermoshentsev, B. Nougmanov, and F. Khalili for useful discussions.

### v3

@@ I Introduction
One of the most well known type of the non-Gaussian states is the Schrödinger cat (SC) one. The “mainstream” method of preparation of these states is based in photon subtraction and conditional measurement, see Refs. [20, 21, 22, 23, 24, 25, 26] and the review [27]. The values of SC state amplitudes up to $\alpha\sim 2$ were demonstrated using this method. Iterative growth of the state of SC was also proposed and demonstrated [28, 26]. The recent estimates [29, 30] show that it allows to prepare SC states with the amplitudes $\alpha\gtrsim 4$ and fidelity exceeding 99%. We would like to mention also the experimental works aimed on increasing the probability of preparation of the SC state [31] and on the remote preparation of the SC state using two-mode squeezing [32]. The influence of optical losses on the SC state was explored experimentally in Ref. [33].
@@
Unfortunately, this method requires a very strong parametric interaction and small linear losses in the signal mode [41]. By now, these conditions were successfully fulfilled only in microwave cryogenic systems with Josephson junctions [42, 43, 44]. Cold-trapped ions [45] and semiconductor microcavities [46, 47, 48] can also be used as possible platforms for various applications of the spontaneous parametric-down conversion.
The joint unitary evolution of the two modes was analyzed in the works [34, 35, 49] numerically, showing that quantum states that resemble SC ones (but not equal to them) can be generated in this process; see more details in Sec. III.
@@ III Unconditional evolution
In Ref. [35], evolution of the partial Husimi quasi-probability function of the signal mode was explored numerically. In Ref. [49], the partial Wigner functions of both the signal and the pump modes were calculated numerically for the small values of $\alpha\sim 2$. However, it has to be emphasized that in both these works, the resulting states were obtained by tracing out the second modes. This procedure gives inherently mixed states that are not so interesting for perspective applications of the non-Gaussian states.
In Fig. 2, the corresponding probability distribution for the photon number in the signal mode is shown. Similar to the “even” SC state
@@ IV Conditional preparation of the Schrödinger cat state
Here we consider the protocol that, in principle, generate the pure quantum state approaching the SC one. Suppose that at some moment of the normalized time $\tau_{\rm opt}$, the photon number $\hat{n}_{p}$ in the pump mode is measured. Suppose also that we are interested only in the events where the value $\hat{n}_{p}=0$ is obtained. The schematic diagram of this procedure is shown in Fig. 3.
Figure 3: The scheme of the conditional preparation of the SC state. Red line – signal mode, blue line – pump mode. SPD – single photon detector. The auxiliary rotation ($\hat{R}^{\dagger}$) and anti-squeezing ($\hat{S}^{\dagger}$) elements are discussed in the end of Sec. IV
@@
Figure 5: Wigner function of the output state $|{\psi_{\rm out\,raw}}\rangle$ at the optimal time $\tau_{\rm opt}$. $\beta=3$
@@
In Fig. 6, deviation of the resulting fidelity from one is plotted as a function of $\beta$. It does not exceed $3\cdot 10^{-3}$ and most probably results from the accumulated errors of the numerical calculations (the dip in the vicinity of $\beta=57$ and several smaller dips in at $\beta=30\text{-}50$, most probably, originate for the peculiarities of the numerical calculation library we used).
If necessary, then the prepared squeezed SC state $|{\psi_{\rm out\,raw}}\rangle$ could be converted to the non-squeezed one $|{\psi_{\rm out}}\rangle$ using the squeezing operation with the negative squeeze factor $-\xi_{\rm prep}$, see Eq. (17), implemented by means of a degenerate optical parametric oscillator in the strong inexhaustible classical pumping mode regime, see e.g. the review [50]. The resulting average photon number can be calculated as follows (see Eq. (30)):
@@
Also if necessary, the tilt of the prepared state can be adjusted using the rotation (that is the phase delay) operation.
V Discussions and conclusions
@@ Acknowledgements.
This work was supported by the Russian Science Foundation (project 20-12-00344) and Theoretical Physics and Mathematics Advancement Foundation “BASIS” (Grant 23-1-1-39-1). The author would like to thank D. Chermoshentsev, B. Nougmanov, and F. Khalili for useful discussions.